Can someone help me solve this?
Answer the following question.
Answer the following question about the system
7x+5y=2
8x-9y=17
If we multiplied the second equation by 5, what number would we have to multiply the first equation by to eliminate the y terms?

Answer by edjones(8007)

8x-9y=17
40x-45y=85 multiplied by 5
.
7x+5y=2 To eliminate y you must multiply by 9 to make the two y's add to zero.
63+45y=18
.
40x-45y=85
63x+45y=18
-----------add
103x = 103 making y disapear, leaving only x.
x=1
.
Ed